# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v12

@@ 2.9 Statistical Bounding
When the underlying distribution is Normal, Formula (2.47) and (2.48) give the distributional bounding leakage $\epsilon(\kappa)$ and the sample bounding leakage $\epsilon(\kappa_{s},N)$ respectively, where $\xi()$ is the Gaussian error function [5]. Figure 5 shows that $\epsilon(\kappa_{s})<\epsilon(\kappa_{s},N)$, and $\lim_{N\rightarrow\infty}\epsilon(\kappa_{s},N)=\epsilon(\kappa_{s})$. It also shows that $\kappa(\kappa_{s},N)<\kappa_{s}$ and $\lim_{N\rightarrow\infty}\kappa(\kappa_{s},N)=\kappa_{s}$. Figure 6 further demonstrates that for smaller $\kappa_{s}$, $\kappa(\kappa_{s},N)$ approaches $\kappa_{s}$ more rapidly as $N$ increases (e.g., $\kappa(2,100)\simeq 2$ vs $\kappa(5,1000)\simeq 5$), but converges to a larger stable bounding leakage (e.g., $\epsilon(2)=4.55\;10^{-2}$ vs $\epsilon(5)=5.73\;10^{-7}$). Figure 6 also indicates that when $N\geq 30$, the difference between $\epsilon(4,N)$ and $\epsilon(5,N)$ is less than $10^{-3}$, suggesting that $\kappa(\kappa_{s},N)$ becomes stable when $\kappa_{s}\geq 4$. Moreover, according to the 5-$\sigma$ rule, $\kappa(5,N)$ in Figure 5 should be used as $\kappa$ in Formula (2.7).
@@ 2.10 Ideal Statistics
  • When the underlying uncertainty distribution is Uniform, Figure 8 shows the same trend, however the stable variances are reached only when $N>10^{5}$. The stable variance $\widehat{\delta^{2}}f$ is computed using Formula (2.18).
@@
In both Figure 7 and 8, the bonding leakage $\epsilon(N)$ decreases as the sample count $N$ increases. The stable resulting variances $\delta^{2}f$ is reached only when $\epsilon(s,N)<10^{-3}$. All functions approach their corresponding ideal variances with exactly the same trend. These observations suggest that bounding leakage is the reason for $\delta^{2}f<\widehat{\delta^{2}}f$ and it is justified to infer confidential interval range $\kappa$ from bounding leakage $\epsilon(N)$. Define $\alpha\equiv\delta^{2}f/\widehat{\delta^{2}}f\in[0,1]$ as the ideal ratio, which quantifies the reliability for $\widehat{\delta^{2}}f$ due to insufficient sample count $N$ for the given underlying distribution. When there are multiple inputs each with its own ideal ratio $\alpha=\zeta(2,\kappa)$, Formula (2.50) demonstrates an example on how to calculate the resulting ideal ratio $\alpha$ for $x\pm y$.
An ideal ratio also applies to the corresponding uncertainty bias. The trends of uncertainty biases $\overline{f}-f$ versus sample count NN are exactly as those in Figure 7 and 8, except that $\overline{x}-x=0$. This similarity is expected because the first-order approximation in both Formula (2.9) and Formula (2.3) contains $\sigma^{2}\zeta(2)$.
@@ 3 Variance Arithmetic
  • stable: To avoid truncation error [13], the value of the last expansion term must be less than $5.73\;10^{-7}$ times of both the result uncertainty and the result absolute value. This rule ensures sufficiently fast convergence in the context of monotonic convergence.
@@ 4.3 Types of Calculations to Verify
An application algorithm computes numerical values from an analytic formula. Through statistical Taylor expansion, variance arithmetic applies directly to analytic problems.
A transformation algorithm produces output data of approximately the same quantity as its input, with the overall information content remaining largely unchanged. For reversible transformations, a unique requirement is to recover every original input for both value and uncertainty after a round-trip transformation that is performing a forward transformation followed by its reverse transformation. The discrete Fourier transform (DFT) is a typical reversible transformation algorithm: it has the same amount of input and output data, and its output can be transformed back into the input using essentially the same process. A test of variance arithmetic using the fast Fourier transform (FFT, which is an implementation of DFT) algorithms is presented in Section 9.
@@ 7.1 Matrix Determinant
Let vector $p_{1},p_{2}\dots p_{n}]_{n}$ denote a permutation of the vector $(1,2\dots n)$ [37]. Let $\mathdollar[p_{1},p_{2}\dots p_{n}]_{n}$ denote the permutation sign of $p_{1},p_{2}\dots p_{n}]_{n} [$37]. Formula (7.1) defines the determinant of a $n$-by-$n$ square matrix $\mathbf{M}$ with the element $x_{i,j},i,j=1\dots n$ [37]. The sub matrix $\mathbf{M_{i,j}}$ at index $(i,j)$ is formed by deleting the row $i$ and column $j$ of $M$, whose determinant is given by Formula (7.2) [37]. For discussion simplicity, sub determinant $|\mathbf{M}|_{i,j}$ in Formula (7.2) contains the permutation sign, which is different from the determinant of the sub matrix $|\mathbf{M}_{i,j}|$ [37] that treats the sub matrix $\mathbf{M_{i,j}}$ as an independent matrix. Formula (7.3) holds for the arbitrary row index $i$ or the arbitrary column index $j$ [37].
@@ 8.1 Moving-Window Linear Regression Algorithm
According to Figure LABEL:fig:_Moving_Linear_Fit_Coeff, in which $H$ takes an example value of 2, the values of $(\alpha_{j},\beta_{j})$ can be derived from the previous values $(\alpha_{j-1},\beta_{j-1})$, allowing Formula (8.3) and (8.4) to be reformulated into the progressive moving-window calculation given by Formula (8.5) and (8.6), respectively [1].
@@ 9.6 Using Quart Sine for Sin/Cos Signals
Figure 36: FFT spectrum of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using the Quart sine function after the forward transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
Figure 37: FFT waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using the the Quart sine function after the reverse transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the time index, while the y-axis represents both the uncertainty and the value error.
Figure 38: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Sin/Cos signal versus FFT order (x-axis) and transformation types (legend) using the Quart sine function.
@@
Using the Quart sine function, for a sine wave with a frequency of $3$, Figure 36 displays the spectrum for the forward transformation, while Figure 37 presents the waveform for the reverse transformation.
  • In the forward transformation, the value errors are slightly smaller than the corresponding result uncertainties, with an error deviation of $0.31$. In the reverse transformation, the error deviation is $0.43$. Both error deviations are less than $1$, confirming a slight overestimation of the sine uncertainty by ULP.
@@
  1. 1.
    As the FFT order increases, the uncertainty mean of the forward transformation grows exponentially faster than that of the reverse transformation. At FFT order $18$, the uncertainty mean of the forward transformation is $10^{3}$ times greater than that of the reverse transformation.
  2. 2.
    At FFT order $18$, Figure 39 shows that the normalized error distribution for the reverse transformation is wider than that of the forward transformation, although the two distributions do not differ by an order-of-magnitude. Therefore, the value error increases at least $10^{2}$ times faster in the forward transformation than in the reverse transformation.
  3. 3.
    In forward transformation, error deviations reach their stable values rapidly once the FFT order $L\geq 4$. In contrast, in the reverse transformation, error deviations stabilize more slowly with increasing FFT order, approaching stability only when $L\geq 15$. This slow convergence of normalized errors in the reverse transformation is attributed to its input data which consists entirely of precise zeros except at the frequency indexes.
Despite these differences, error deviations are between $0.2$ and $0.5$, suggesting that in variance arithmetic, uncertainties track value errors effectively in all cases, achieving proper coverage.
@@ 9.7 Using Library Sine for Sin/Cos Signals
Figure 40: FFT spectrum of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the forward transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
Figure 41: FFT waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the reverse transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the time index, while the y-axis represents both the uncertainty and the value error.
@@
Using the Library sine function, for a sine wave with a frequency of $3$, Figure 40 displays the spectrum for the forward transformation, while Figure 41 presents the waveform for the reverse transformation:
  • The uncertainties are identical to the corresponding values obtained using the Quart sine function, since the input uncertainties are the same for both sine functions. This identical relationship persists across all FFT orders when comparing the corresponding uncertainty means in Figure 38 and 42.
@@
  • Due to numerical errors in the Library sine function, error deviations reach their stable values much quicker with FFT order in Figure 42 that that in Figure 38.
Using the Library sine function, variance arithmetic also achieves proper coverage for sine and cosine signals.
@@ 9.8 Numerical Error Resonance Using Library Sine
In the reverse transformation, value errors exhibit a clear trend of increasing with the time index, as shown in Figure 44. These large value errors appear systematic rather than random and visually resemble a resonant pattern. Similar increases are observed at in other frequencies and FFT orders, as well as in computational results obtained using mathematical libraries such as SciPy. In the forward transformation, although there is no visible resonant pattern in FFT spectrum, the intensity of value errors increase with FFT order. Figure 44 and 45 demonstrate that the error deviations increase with sine or cosine frequency. In contrast, this increase is completely absent when using the Quart sine function, as seen in Figure 36 and 37. Figure 34 indicates that the numerical errors using the Library sine function increase with a periodicity of $\pi$, which may resonate with a signal whose periodicity of an integer multiply of $\pi$, producing the resonant pattern in Figure 40 and 41. At higher frequency, the resonant beats between the signal and the numerical errors in the Library sine function become stronger. To suppress this numerical error resonances, an input noise level of approximately $10^{-14}$ input noise must be added to the sine or cosine signals.
@@ 9.9 Using Quart Sine for Linear Signal
Figure 46 shows that using the Quart sine function, the result uncertainties can track the value errors of Linear signals with proper coverage. Because the input to the reverse transformation no longer consists predominantly of precise zeros, the output uncertainty increases much more rapidly with FFT order than its counterpart in Figure 38. This increase enables the proper coverage in Figure 46, resulting in error deviations similar to those Figure 38. The normalized error histogram of the reverse transformation in Figure 49 is narrower than that of Figure 43, confirming Figure 46. Using the Quart sine function, variance arithmetic can provide proper coverage for Sin signals, Cos signals, and Linear signal.

### v13

@@ 2.9 Statistical Bounding
When the underlying distribution is Normal, Formula (2.47) and (2.48) give the distributional bounding leakage $\epsilon(\kappa)$ and the sample bounding leakage $\epsilon(\kappa_{s},N)$ respectively, where $\xi()$ is the Normal error function [4]. Figure 5 shows that $\epsilon(\kappa_{s})<\epsilon(\kappa_{s},N)$, and $\lim_{N\rightarrow\infty}\epsilon(\kappa_{s},N)=\epsilon(\kappa_{s})$. It also shows that $\kappa(\kappa_{s},N)<\kappa_{s}$ and $\lim_{N\rightarrow\infty}\kappa(\kappa_{s},N)=\kappa_{s}$. Figure 6 further demonstrates that for smaller $\kappa_{s}$, $\kappa(\kappa_{s},N)$ approaches $\kappa_{s}$ more rapidly as $N$ increases (e.g., $\kappa(2,100)\simeq 2$ vs $\kappa(5,1000)\simeq 5$), but converges to a larger stable bounding leakage (e.g., $\epsilon(2)=4.55\;10^{-2}$ vs $\epsilon(5)=5.73\;10^{-7}$). Figure 6 also indicates that when $N\geq 30$, the difference between $\epsilon(4,N)$ and $\epsilon(5,N)$ is less than $10^{-3}$, suggesting that $\kappa(\kappa_{s},N)$ becomes stable when $\kappa_{s}\geq 4$. Moreover, according to the 5-$\sigma$ rule, $\kappa(5,N)$ in Figure 5 should be used as $\kappa$ in Formula (2.7).
$\kappa_{s}=\sqrt{3}$ and $\kappa_{s}=5$ are defined as the ideal bounding ranges for the uniform and Gaussian distributions respectively.
@@ 2.10 Ideal Statistics
  • When the underlying uncertainty distribution is Uniform, Figure 8 shows the same trend, however the stable variances are reached only when $N>10^{4}$. The stable variance $\widehat{\delta^{2}}f$ is computed using Formula (2.18).
@@
In both Figure 7 and 8, the bonding leakage $\epsilon(N)$ decreases as the sample count $N$ increases. The stable resulting variances $\delta^{2}f$ is reached only when $\epsilon(s,N)<10^{-3}$. All functions approach their corresponding ideal variances with exactly the same trend. These observations suggest that bounding leakage is the reason for $\delta^{2}f<\widehat{\delta^{2}}f$ and it is justified to infer confidential interval range $\kappa$ from bounding leakage $\epsilon(N)$. Define $\alpha\equiv\delta^{2}f/\widehat{\delta^{2}}f\in[0,1]$ as the ideal ratio, which quantifies the reliability for $\widehat{\delta^{2}}f$ due to insufficient sample count $N$ for the given underlying distribution. Define the ideal ratio to be zero when the sample count $N$ is $1$. When there are multiple inputs each with its own ideal ratio $\alpha=\zeta(2,\kappa)$, Formula (2.50) demonstrates an example on how to calculate the resulting ideal ratio $\alpha$ for $x\pm y$.
An ideal ratio also applies to the corresponding uncertainty bias. The trends of uncertainty biases $\overline{f}-f$ versus sample count $N$ are exactly as those in Figure 7 and 8, except that $\overline{x}-x=0$. This similarity is expected because the first-order approximation in both Formula (2.9) and Formula (2.3) contains $\sigma^{2}\zeta(2)$.
Statistical Taylor expansion outputs ideal uncertainty, ideal uncertainty bias, and ideal ratio for the chosen ideal bounding ranges, in which the ideal ratio quantifies the reliability of the ideal uncertainty and ideal uncertainty bias.
@@ 3 Variance Arithmetic
  • stable: To avoid truncation error [12], the value of the last expansion term must be less than $5.73\;10^{-7}$ times of both the result uncertainty and the result absolute value, in which $5.73\;10^{-7}$ is the bounding leakage for Gaussian distribution with bounding range $\kappa=5$. This rule ensures sufficiently fast convergence in the context of monotonic convergence.
@@ 4.3 Types of Calculations to Verify
An application algorithm computes numerical values from an analytic formula. Through statistical Taylor expansion, variance arithmetic applies directly to analytic problems. For example, for the catastrophic cancellation example of Formula (1.2), Formula (4.3) shows that variance arithmetic bound the rounding error of $1$ with uncertainty $1$:
$64919224$ $\times 205117922-159018721\times 83739041=$
$13316075197586562.-13316075197586560.\pm 1=2.\pm 1;$ (4.1)
A transformation algorithm produces output data of approximately the same quantity as its input, with the overall information content remaining largely unchanged. For reversible transformations, a unique requirement is to recover every original input for both value and uncertainty after a roundtrip transformation that is performing a forward transformation followed by its reverse transformation. The discrete Fourier transform (DFT) is a typical reversible transformation algorithm: it has the same amount of input and output data, and its output can be transformed back into the input using essentially the same process. A test of variance arithmetic using the fast Fourier transform (FFT, which is an implementation of DFT) algorithms is presented in Section 9.
@@ 7.1 Matrix Determinant
Let vector $p_{1},p_{2}\dots p_{n}]_{n}$ denote a permutation of the vector $(1,2\dots n)$ [14]. Let $\mathdollar[p_{1},p_{2}\dots p_{n}]_{n}$ denote the permutation sign of $p_{1},p_{2}\dots p_{n}]_{n} [$14]. Formula (7.1) defines the determinant of a $n$-by-$n$ square matrix $\mathbf{M}$ with the element $x_{i,j},i,j=1\dots n$ [14]. The sub matrix $\mathbf{M_{i,j}}$ at index $(i,j)$ is formed by deleting the row $i$ and column $j$ of $M$, whose determinant is given by Formula (7.2) [14]. For discussion simplicity, sub determinant $|\mathbf{M}|_{i,j}$ in Formula (7.2) contains the permutation sign, which is different from the determinant of the sub matrix $|\mathbf{M}_{i,j}|$ [14] that treats the sub matrix $\mathbf{M_{i,j}}$ as an independent matrix. Formula (7.3) holds for the arbitrary row index $i$ or the arbitrary column index $j$ [14].
@@ 8.1 Moving-Window Linear Regression Algorithm
The values of $(\alpha_{j},\beta_{j})$ can be derived from the previous values $(\alpha_{j-1},\beta_{j-1})$, allowing Formula (8.3) and (8.4) to be reformulated into the progressive moving-window calculation given by Formula (8.5) and (8.6), respectively [29].
@@ 9.6 Using Quart Sine for Sin/Cos Signals
Figure 36: FFT value error specturm of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using the Quart sine function after the forward transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
Figure 37: FFT value error waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using the the Quart sine function after the reverse transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the time index, while the y-axis represents both the uncertainty and the value error.
Figure 38: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Sin/Cos signal versus FFT order (x-axis) and transformation types (legend) using the Quart sine function. The variance arithmetic is implemented using python.
@@
Using the Quart sine function, for a sine wave with a frequency of $3$, Figure 36 displays the value error spectrum for the forward transformation, while Figure 37 presents the value error waveform for the reverse transformation.
  • In the forward transformation, the value errors are slightly smaller than the corresponding result uncertainties, with an error deviation of $0.60$. In the reverse transformation, the error deviation is $0.55$. Both error deviations are less than $1$, confirming a slight overestimation of the sine uncertainty by ULP.
@@
  • The value error waveform for the roundtrip transformation is qualitatively similar to that in Figure 37 with a similar error deviation of $0.55$.
@@
  • In the forward transformation, error deviations reach their stable values around $0.65$ rapidly once the FFT order $L\geq 6$.
  • In contrast, in the reverse transformation, error deviations stabilize more slowly with increasing FFT order, approaching their stable values around $0.90$ only when $L\geq 12$. This slow convergence of normalized errors in the reverse transformation is attributed to its input data which consists entirely of precise zeros except at the frequency indexes.
  • At FFT order $18$, Figure 39 shows that the normalized error distribution for the reverse transformation is wider than that of the forward transformation, confirming the larger error deviation.
The value error increases at least $10^{2}$ times faster in the forward transformation than in the reverse transformation. Despite these differences, error deviations are $0.65$ and $0.90$, suggesting that in variance arithmetic, uncertainties track value errors effectively in all cases, achieving proper coverage.
9.7 Sensitivity to Quart sine functions
Figure 40: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Sin/Cos signal versus FFT order (x-axis) and transformation types (legend) using the Quart sine function. The variance arithmetic is implemented using C++.
Figure 41: Histograms of normalized errors between the Quart sine functions of C++, Java, and Python in the unit of least significant bits.
The reverse transformation of Sin/Cos signals is very sensitive to the underlying Quart sine function. If variance arithmetic is implemented using C++ instead of Python, Figure 40 shows that the error deviations of the reverse transform quickly reaches the stable values around $0.89$, while the forward transformation remains similar to its counterpart in Figure 38. Figure 40 also shows identical uncertainty means to their counterparts in Figure 38. Figure 41 shows that the difference of the Quart sine functions between Python and C++ is only about $1\%$ LSBs. Figure 41 also reveals that the difference of the Quart sine functions between Java and C++ is only about $0.1\%$ LSBs. Indeed, the results between Java and C++ are much more similar. Because the Python implementation gives smaller error deviations, its Quart sine function has less numerical errors on average. For discussion simplicity, the rest results are only based on the python implementation.
@@ 9.8 Using Library Sine for Sin/Cos Signals
Figure 42: FFT value error specturm of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the forward transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
Figure 43: FFT value error waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the reverse transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the time index, while the y-axis represents both the uncertainty and the value error.
@@
Using the Library sine function, for a sine wave with a frequency of $3$, Figure 42 displays the value error spectrum for the forward transformation, while Figure 43 presents the vale error waveform for the reverse transformation:
  • The uncertainties are identical to the corresponding values obtained using the Quart sine function, since the input uncertainties are the same for both sine functions. This identical relationship persists across all FFT orders when comparing the corresponding uncertainty means in Figure 38 and $4$.
@@
  • Due to numerical errors in the Library sine function, error deviations reach their stable values much quicker with FFT order in Figure 44 that these in Figure 38.
  • The stable error deviations using the Library sine function are much larger than their counterparts using Quart library functions: $10.1>0.65$ for the forward transformation, $6.3>0.90$ for the reverse transformation, and $1.47>0.94$ for the roundtrip transformation.
Using the Library sine function, variance arithmetic barely achieves proper coverage for sine and cosine signals.
@@ 9.9 Numerical Error Resonance Using Library Sine
Figure 46: FFT value error waveform of $\cos(6\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the reverse transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
Figure 47: FFT value error waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the roundtrip transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
@@
In the reverse transformation, value errors exhibit a clear trend of increasing with the time index, as shown in Figure 48. These large value errors appear systematic rather than random and visually resemble a resonant pattern. Similar increases are observed at in other frequencies and FFT orders, as well as in computational results obtained using mathematical libraries such as SciPy. For example, Figure LABEL:fig:_FFT_Cost_Clean_6_6_Wave_Lib shows that the FFT value error waveform of $\cos(6\frac{2\pi}{2^{6}}j)$ has much stronger resonant pattern than that of $\sin(3\frac{2\pi}{2^{6}}j)$ in Figure 43. In the forward transformation, although there is no visible resonant pattern in FFT value error specturm, the intensity of value errors increase with FFT order. In contrast, such resonance is absent for the roundtrip transformation as shown in Figure 47. Also, it is completely absent when using the Quart sine function, as seen in Figure 36 and 37. Figure 48 and Figure 49 demonstrate that the error deviations increase with sine or cosine frequency. Figure 34 indicates that the numerical errors using the Library sine function increase with a periodicity of $\pi$, which may resonate with a signal whose periodicity of an integer multiply of $\pi$, producing the resonant pattern in Figure 42, 43, and 46. At higher frequency, the resonant beats between the signal and the numerical errors in the Library sine function become stronger. To suppress this numerical error resonances, an input noise level of approximately $10^{-14}$ input noise must be added to the sine or cosine signals.
@@ 9.10 Using Quart Sine for Linear Signal
Figure 50 shows that using the Quart sine function, the result uncertainties can track the value errors of Linear signals with proper coverage. Because the input to the reverse transformation no longer consists predominantly of precise zeros, the output uncertainty increases more rapidly with FFT order than its counterpart in Figure 38. Such increase results in larger error deviations: At FFT order $18$, $0.89>0.65$ for the forward transformation, $1.08>0.90$ for the reverse transformation, $1.67>0.90$ for the roundtrip transformation, and wider distributions of the normalized value errors in Figure 51 versus the corresponding histograms in Figure 39. Using the Quart sine function, variance arithmetic can provide proper coverage for Sin signals, Cos signals, and Linear signal.